OpenStudy (anonymous):

hey find the points of inflection on f(x)=6x^4-x^2 thanksss oh and i got plus or minus 1/6

OpenStudy (anonymous):

was that not correct?

OpenStudy (anonymous):

f'(x) = 24x^3 - 2x f''(x) = 72x^2 - 2 set f''(x) = 0...you get 2(36x^2 - 1) = 0, or x = +1/6 and -1/6 Now, set up an interval chart to see where concavity, if any, exists.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

do you plug in the number to the first der. ?

OpenStudy (anonymous):

to check for concavity